OpenStudy (anonymous):

The probability that a professional baseball player will get a hit is 1/3. Calculate the exact probability that he will get ATLEAST 3 hits in 5 attempts. And explain how you do this, please?

OpenStudy (anonymous):

5/15

OpenStudy (anonymous):

.3333333333->

OpenStudy (zarkon):

assuming independence... \[\sum_{k=3}^{5}{5 \choose k}\left(\frac{1}{3}\right)^{k}\left(1-\frac{1}{3}\right)^{5-k}\]